OpenStudy (anonymous):

:) How do you write the ratio as a fraction in lowest terms? Please explain. Example 1: 3 to 5

OpenStudy (anonymous):

If I'm correct all it is would be 3/5. Simple. @Gin_Ichimaru I would like Gin to check to make sure

OpenStudy (anonymous):

Or would it be .6?

OpenStudy (anonymous):

I do not understand the question There are three ways to write a ratio. For Example: 3:5 3/5 3 to 5

OpenStudy (anonymous):

There are other forms... I have lot of these to do. Here are some more examples... 1. 3 to 5 2. 2 to 6 3. 8 to 24 4. 6:4 5. 6:8

OpenStudy (anonymous):

Those are just different numbers the forms are all the same 3.

OpenStudy (anonymous):

3/5 1/4 1/4 3/2 3/4

OpenStudy (anonymous):

If I understood the question correctly, that should be what youre looking for. Need anymore help?

OpenStudy (anonymous):

3/5 2/6=1/3 8/24=1/3 6/4=3/2 6/8=3/4 Do you understand

OpenStudy (anonymous):

So I'm basically dividing them by the same number they have in common?

OpenStudy (anonymous):

Sorry I wrote 1/4 it was 1/3 lol

OpenStudy (anonymous):

Make them fractions.

OpenStudy (anonymous):

Ya. So I'm basically simplifying them in fraction form... correct?

OpenStudy (anonymous):

Yep :) You got it!

OpenStudy (anonymous):

Thanks :)